Commander Report – April 3rd
Crew 227 Commander Report 3st April 2022
Name of person filing report: Cyril Wain
Thank you for reading us!
Today, everyone wakes up early and with great motivation since it was the big day and this for several reasons. Indeed, this morning a part of crew went on EVA to discover an unexplored canyon, we ate pancakes at breakfast, and we had the visit of a journalist called Emily.
Regarding the EVA, the team was constituted of Julie, Jean, Cheyenne, Sirga and I. At 9h00, we were ready to start our exploration. The rule was simple: never stopped and walked fast to go as far as we can. And the thing is that it worked quite well! We went super far and achieved a wonderful canyon!
I used the drone to make some videos of it. Then, I used another software to start doing the 3D model of the canyon. However, the data acquired for the 3D model overwrite the videos and cancelled them. I discovered that only when we came back at the hab and that I wanted to show the videos to everyone… and the worst in the story is that, while doing the 3d modelling of the canyon, I lost the GPS connection so there are no GPS coordinates in the metadata of my pictures which makes the 3D modelling impossible to proceed. We can definitely call that a big fail. It is not a big issue since, now, I am aware of that and I will not make that error again. But it is really disappointing since we will not go to that canyon anymore and will not be able to perform its 3d modelling. Tomorrow, I would like to perform the 3D modelling of the MDRS.
Hopefully, we have our own memory and we will always remember this amazing canyon! Also, we can count on the pictures that Julie took.
During that time, Julien, Audrey and Ignacio met Emily. She is coming from Boston but now is living in Moab. They present the whole station to her and explained her their respective experiment.
At 12h30, we ate all together. We ate the rest of yesterday’s dinner. During this lunch, we spoke a lot with Emily about the reason we are here doing this atypical experiment. She also spoke about her own life. It was very nice.
Unfortunately, between 14h00 and 16h00, Audrey was quite sick. Hopefully, with our great HSO Julie, Audrey was back on foot quite fast.
At that time, the EVA crew, explained their personal experiment to Emily. She was happy of her visit and left the station around 15h30.
Then, Julien, Audrey, Sirga and Jean did a match for the experiment of Julie. For her first match, Audrey did a great job and won against Sirga!
At the same time, Ignacio, Cheyenne and I worked on our side.
Now, as every day, we are all together in the kitchen while doing our reports for the capcom windows.
Yesterday night, we ate some potatoes with non-mixed crickets. It was difficult to eat for the whole team. The taste was ok but the visual and texture was quite bad (for us). It was a bad point for the moral of the crew. After the dinner, Julie did a surprise match. It was Sirga against Ignacio. Ignacio won and bring one more point to the engineering team!
However, the moral of the team is still very good! Tonight, we are going to eat chicken with rice and curry and moreover without any insects!! After, we plan to play some cards game.
Regarding the station, everything is nominal.
Operations Report – April 3rd
Crew 227 Operations Report 03-04-2022
Name of person filing report: Julien Meert
Non-nominal systems: Nothing to report
Notes on non-nominal systems: Nothing to report
Spirit rover used: Yes
Ending charge: 84
Currently charging : We plug it after the EVA
Opportunity rover used: Yes
Ending charge: 85
Currently charging: Plug in
Curiosity rover used: Yes
Beginning charge: 100%
Ending charge: 86
Currently charging: plug in
Perseverance rover used: Not use today
General notes and comments: Nothing to report
Summary of Hab operations : Ignacio pored his coffee on his computer but we manage to repair it
Water (static tank): 312 gallons
Water (loft tank): 25 gallons
Water Meter: 157232,2 units
Static to Loft Pump used : 38 gallons
Static tank pipe heater (on or off): Off
Static tank heater Off
Toilet tank emptied: Not today
Summary of internet: Stable
Summary of suits and radios: Nothing new
Summary of GreenHab operations: Nothing new
WATER USE: 0 gallons
Supplemental light: Off
Summary of ScienceDome operations: We will let the incubator ON during the night for an experimentation
Dual split: Off
Summary of RAM operations: Nothing to report
Summary of any observatory issues: Nothing to report
Summary of health and safety issues: Nothing to report
Questions, concerns and requests to Mission Support: Nothing but saying thank you for your work !
Journalist Report – April 3rd
Crew 227 Journalist Report 3 April 2022
Sol : 7
Name of the person filling the report : Sirga Drouet
Today we got up early but slept late, for some of us, after a surprise game until 11 pm. Team mechanics against team the "Others" Ignacio against me the astronauts will be able to return on earth, good news even if Ignacio remains the best of the apprentice surgeons although we all evolve!
This morning breakfast of quality, good pancakes with syrup, bleuberries and even chocolate!
We ate in two seconds our good meal because a very precise goal and a mission para commando to realize was of setting.
Return to Tharsis canyon, to go as far as possible in three hours by counting the way in rover. We clearly made the fastest preparation of EVA since we are here. At 9:16 am sharp we were in front of the tricky canyon entrance. Euphorically we set off, plunging into what would turn into a wall of rock. The walk was fast, I admit it was clearly military for me with my little legs. We went two and a half times faster than the day before ! The team was composed of Julie, Jean, Cyril, me and Cheyenne. The easiest way to advance was to go along the ridge to have a better view.
We were higher and higher, we felt so ridiculous in front of these strata which are and will be there before and maybe after the humanity. A funny feeling, it pushes us to question ourselves on our place in this world and on our so small trace in the rocks in front of us. We stayed there a moment to contemplate from the top of the 40 meters the river dry in the bottom of the valley. Some shots of drone and in spite of the fact that we were in 2/3 of our objective it did not remain us enough time and the half-turn arrived.
To finish this canyon we would have to run at each negative drop, make a quick hello at the origin and run back without forgetting our 10 Kilograms of space suit. We are proud of what we have accomplished so far and maybe the origin of our goal will remain forever blank. An ideal to be reached like the utopia of making humanity live on Mars. But I am sure that the echoes of our hopes will plausibly solve terrestrial problems useful to our future.
The secret in this adventure is to not give up and despite the many failures to stay united and face the challenge.
The return journey was less sportive but the sun was shining and the thirst was felt. On the way back we got the antibiotics from Audrey who was exposed.
Once back, the routine : The decompression chamber, undress and eat. Today no insects of the day to take a break, everyone seems to be in heaven ahaha.
The afternoon activities started and Audrey’s first surgical match took place, against me, she did very well for a first.
Then a little manipulation of Audrey’s bacteria and Cheyenne’s plants are growing wonderfully !
During this time it was the church in the Hab to pray of Saint-Computer because an accident happened… We almost lost the computer of Ignacio because of an attack of coffee. We were three reconverted in monk while Father Cyril with his hands of electrical engineer handled the multimeter to try to find the cause of the murder. The machine was making big beeps, which is very worrying in the microscopic computer world… Finally, our prayers were partly heard and the only collateral damage was to remove the battery.
A good day passed, tonight card tournaments are on the program to blow.
Astronomy Report – April 3rd
Name: Ignacio Sánchez Casla
Sky Conditions: High altitude clouds
Wind Conditions: No wind
MUSK SOLAR OBSERVATORY
Observation Start Time: 14:00
Observation End Time: 15:00
Summary: Tried to use the solar observatory this afternoon but due to the very unstable conditions created by the clouds blocking then unblocking the Sun’s view I couldn’t take any meaningful pictures.
Objects Viewed: Sun
Problems Encountered: None
MDRS ROBOTIC OBSERVATORY
Robotic Telescope Requested: MLC-RCOS16
Objects to be imaged this evening:
– Messier 51
– Messier 64
– Messier 57
– Messier 101
– NGC 7023
The usual suspects
Images submited with this report:
None. Had a problem with my computer this afternoon – poured some coffee on it and he didn’t really liked it, he’s fine now – so I will process last night pictures this evening.
MDRS-WF seems not to be looking in right direction. I know that it is not entirely operational and that you are monitoring the situation so no worries on my side. I won’t use it tonight tough.
Some images failed on the MLC-RCOS16 but most of them were fine.
I have no credits left on the New Mexico Observatories. Since I will only be using the MLC-RCOS16, can I get my credits transfered?
EVA Report – April 3rd
Crew 227 EVA Report 3rd April 2022
Name of person filing report: Jean Jacobs
Purpose of EVA: Further exploration of the Tank wash starting at Cow Dung Road
Start time: 9:00
End time: 12:00
Narrative: The team drove up 500m North of reservoir Dam on Cow Dung Road and parked the rovers in front of the beginning of Tank wash as did yesterday’s team. Then they entered the wash and directly went as far as they could. They were able to reach 425500 N; 520750 E a pillar rock in the middle of the canyon about 2.5 km within. We fired up the drone to take pictures for future 3D modeling. The canyon layers seemed like sheets of rocks making it look like a book on the side. Unfortunately, the picture taken lacked the GPS coordinate to make a 3D map of the canyon making it impossible to map the canyon.
Despite a cloudy morning, we did not have any rain.
Destination: Tank wash North East of Cow Dung road
Coordinates: 425500 N; 520750 E
Participants: Jean, Cyril, Sirga, Julie, Cheyenne
Roads and routes per MDRS map: Cow Dung Road 0110
Mode of travel: Rovers Spirit, Opportunity and Curiosity
Sol Summary – April 3rd
Crew 227 Sol Summary Report 03/04/22
Summary Title: A stranger puts the crew under the microscope plus technical difficulties slow down our projects.
Author’s name: Jean Jacobs (XO)
Mission Status: The morning EVA proved to be hard but the reward was worth it: a newly discovered landscape. In the afternoon, the different crew members showed the journalist the base around while working on their experiment.
Technical problems arose during the afternoon with clouds in the sky hiding the sun, computer problems needed a fix and pictures of the EVA lacking GPS coordinates.
Sol activity summary:
09:00 Arrival of the journalist
12:00 Return from the EVA
14:00 Scientific work
– Plants watering and measurement
– Sun picture taken from the Astronomer telescope
– Measurement of Julie’s experiment
– Failed 3D mapping of the newly discovered canyon
– Failure to take sun pictures.
– Ignacio computer repair
18:00 Report redaction
Anomalies in work: None
Weather: Cloudy day, low winds.
Crew Physical Status: A crew member felt dizzy (food intolerance), HSO is monitoring the situation
EVA: Further exploration of the canyon North East of the HAB
Reports to be filed:
– Crew commander
– Operation report
– GreenHab officer
– Picture of the day
– Some pictures
– EVA request
– EVA summary
Support Requested: None at this time
You must be logged in to post a comment.